Hi, please help me solve this equation on radical expressions:
sqrt 14u-12 = sqrt 10u-8
Thanks for any help!

Answer by checkley71(8403)

sqrt(14u-12)=sqrt(10u-8)
square each term:
14u-12=10u-8
14u-10u=-8+12
4u=4
u=1 answer.
proof
sqrt(14*1-12)=sqrt(10*1-8)
sqrt(14-12)=sqrt(10-8)
sqrt(2)=sqrt(2)